# Dusit Niyato, , Bo Ai

### v1

@@ I Introduction
To achieve demanding requirements for wireless communication networks, multiple-input multiple-output (MIMO) technology has been regarded as the promising enabler since fourth-generation (4G) wireless communication networks [1, 2, 3]. Notably, MIMO technology has gathered a big success in the fifth-generation (5G) wireless communication networks since it can deliver higher spectral efficiency in bands that feature advantageous propagation conditions [4, 5]. However, with the further development of wireless communication, sixth-generation (6G) networks are expected to meet substantially higher performance requirements [6, 7, 8] and support a diverse array of emerging application scenarios [9, 10, 11]. To fulfill the anticipated performance enhancements, next-generation MIMO technologies have garnered significant research attention, represented by cell-free massive MIMO (CF mMIMO) [12, 13, 14], reconfigurable intelligent surfaces (RIS) [15, 16, 17], and extremely large-scale MIMO (XL-MIMO) [18, 19, 20]. Among these technologies, CF mMIMO attracted considerable research interest due to its innovative network topology and its ability to provide uniform coverage.
@@
  • For the centralized processing scheme, we study the OBE combining scheme, called centralized OBE (C-OBE), which can maximize the SE among all BE-structured combining schemes. Furthermore, we present a novel closed-form computation of the C-OBE combining scheme.
@@ II Fundamentals of CF mMIMO Networks
In this paper, we study a CF mMIMO network, where $M$ APs with $N$ antennas each serve $K$ single-antenna UEs. All APs, which are connected via fronthaul to a CPU, and UEs are arbitrarily located in a wide coverage region. We define the channel response between AP $m$ and UE $k$ as $\mathbf{g}_{mk}\in\mathbb{C}^{N}$ which remain constant in time-frequency coherence blocks with $\tau_{c}$ channel uses [5]. Note that the uplink pilot transmission and uplink data transmission are investigated. Thus, in each coherence time-frequency block with $\tau_{c}$ channel uses, $\tau_{p}$ and $\tau_{c}-\tau_{p}$ channel uses are reserved for pilots and data, respectively.
@@ II-A Channel Model
We consider the spatially correlated Rician fading channel model, where both the LoS component and NLoS component are modelled. The channel between AP $m$ and UE $k$ can be modelled as $\mathbf{g}_{mk}=\bar{\mathbf{g}}_{mk}e^{j\theta_{mk}}+\check{\mathbf{g}}_{mk},$ where $\bar{\mathbf{g}}_{mk}\in\mathbb{C}^{N}$ is the semi-deterministic LoS component with $\theta_{mk}\sim\mathcal{U}[-\pi,\pi]$ being the random phase-shift between AP $m$ and UE $k$ as in [26] and $\check{\mathbf{g}}_{mk}\sim\mathcal{N}_{\mathbb{C}}\left(\mathbf{0},\check{\mathbf{R}}_{mk}\right)$ is the stochastic NLoS component. $\check{\mathbf{R}}_{mk}\in\mathbb{C}^{N\times N}$ denotes the spatial correlation matrix with $\beta_{mk}^{\mathrm{NLOS}}={\mathrm{tr}(\check{\mathbf{R}}_{mk})}/{N}$ being the large-scale fading coefficient for the NLoS component between AP $m$ and UE $k$. Meanwhile, we assume that $\check{\mathbf{g}}_{mk}$ for every AP-UE pair are independent and channel realizations of $\check{\mathbf{g}}_{mk}$ in different coherence blocks are independent and identically distributed.
@@ II-B Channel Estimation
We use a set of $\tau_{p}$ mutually orthogonal pilot signals, $\boldsymbol{\phi}_{1},\dots,\boldsymbol{\phi}_{\tau_{p}}$, during the channel estimation with $\left\|\boldsymbol{\phi}_{t}\right\|^{2}=\tau_{p}$. Let $t_{k}\in\{1,\dots,\tau_{p}\}$ denote the index of the pilot signal sent by UE $k$. Thus, we denote the pilot signal sent by UE $k$ as $\boldsymbol{\phi}_{t_{k}}$. Due to the limited coherence block size, more than one UE would use the same pilot signal, that is $K>\tau_{p}$. We denote the subset of UEs, applying the same pilot signal as UE $k$, as $\mathcal{P}_{k}$ including UE $k$. All UEs transmit their respective allocated pilot signals and the received pilot signal at AP $m$ can be denoted as $\mathbf{Y}_{m}^{p}=\sum_{m=1}^{M}{\sqrt{p_{k}}\mathbf{g}_{mk}\boldsymbol{\phi}_{t_{k}}^{T}}+\mathbf{N}_{m}^{p}\in\mathbb{C}^{N\times\tau_{p}}$, where $p_{k}$ is the transmit power of UE $k$ and $\mathbf{N}_{m}^{p}\in\mathbb{C}^{N\times\tau_{p}}$ is the additive noise with independent $\mathcal{N}_{\mathbb{C}}\left(0,\sigma^{2}\right)$ entries with $\sigma^{2}$ being the noise power. Based on the fundamentals in [26, 5], we derive the phase-aware MMSE estimate [^1] We assume that the phase-shifts remain the same over many coherence blocks in the frequency domain and only evolve randomly over time. Thus, each AP is assumed to be aware to the phase-shifts. of $\mathbf{g}_{mk}$ as
@@
$\hat{\mathbf{g}}_{mk}=\bar{\mathbf{g}}_{mk}e^{j\theta_{mk}}+\sqrt{p_{k}}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\left(\mathbf{y}_{mk}^{p}-\bar{\mathbf{y}}_{mk}^{p}\right),$ (1)
@@
where $\mathbf{y}_{mk}^{p}=\mathbf{Y}_{m}^{p}\boldsymbol{\phi}_{t_{k}}^{*}=\sum_{l\in\mathcal{P}_{k}}{\sqrt{p_{l}}\tau_{p}\mathbf{g}_{ml}}+\mathbf{n}_{mk}^{p}$, $\bar{\mathbf{y}}_{mk}^{p}=\sum_{l\in\mathcal{P}_{k}}{\sqrt{p_{l}}\tau_{p}\bar{\mathbf{g}}_{ml}}e^{j\theta_{ml}}$, $\mathbf{n}_{mk}^{p}=\mathbf{N}_{m}^{p}\boldsymbol{\phi}_{k}^{*}\sim\mathcal{N}_{\mathbb{C}}(\mathbf{0},\tau_{p}\sigma^{2}\mathbf{I}_{N})$, and $\mathbf{\Psi}_{mk}={\mathbb{E}\{(\mathbf{y}_{mk}^{p}-\bar{\mathbf{y}}_{mk}^{p})\left(\mathbf{y}_{mk}^{p}-\bar{\mathbf{y}}_{mk}^{p}\right)^{H}\}}/{\tau_{p}}=\sum_{l\in\mathcal{P}_{k}}{p_{l}\tau_{p}\check{\mathbf{R}}_{ml}}+\sigma^{2}\mathbf{I}_{N}.$ The estimation error $\tilde{\mathbf{g}}_{mk}=\mathbf{g}_{mk}-\hat{\mathbf{g}}_{mk}$ and estimate $\hat{\mathbf{g}}_{mk}$ are independent random variables, where $\mathbb{E}\{\hat{\mathbf{g}}_{mk}|\theta_{mk}\}=\bar{\mathbf{g}}_{mk}e^{j\theta_{mk}}$, $\mathrm{Cov}\{\hat{\mathbf{g}}_{mk}|\theta_{mk}\}=\hat{\mathbf{R}}_{mk}=p_{k}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk}$, $\mathbb{E}\{\tilde{\mathbf{g}}_{mk}\}=\mathbf{0}$, and $\mathrm{Cov}\{\tilde{\mathbf{g}}_{mk}\}=\mathbf{C}_{mk}=\check{\mathbf{R}}_{mk}-p_{k}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk}$, respectively.
@@ III-A Centralized Processing
$\mathbf{g}_{k}=\left[\mathbf{g}_{1k}^{T},\dots,\mathbf{g}_{Mk}^{T}\right]^{T}=\bar{\mathbf{g}}_{k}+\check{\mathbf{g}}_{k}\in\mathbb{C}^{MN\times 1}$ (3)
where $\bar{\mathbf{g}}_{k}=[\bar{\mathbf{g}}_{1k}^{T}e^{j\theta_{1k}},\dots,\bar{\mathbf{g}}_{Mk}^{T}e^{j\theta_{Mk}}]^{T}$ and $\check{\mathbf{g}}_{k}=[\check{\mathbf{g}}_{1k}^{T},\dots,\check{\mathbf{g}}_{Mk}^{T}]^{T}$ denote the collective LoS and NLoS components, respectively. Relying on the independent characteristic for different AP-UE pairs, we have $\check{\mathbf{R}}_{k}=\mathbb{E}\{\check{\mathbf{g}}_{k}\check{\mathbf{g}}_{k}^{H}\}=\mathrm{diag}(\check{\mathbf{R}}_{1k},\dots,\check{\mathbf{R}}_{Mk})\in\mathbb{C}^{MN\times MN}$. Based on the received pilot signals from all APs $\{\mathbf{Y}_{m}^{p}:m=1,\dots,M\}$, the CPU can obtain the MMSE estimate of $\mathbf{g}_{k}$ as $\hat{\mathbf{g}}_{k}=\bar{\mathbf{g}}_{k}+\sqrt{p_{k}}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}(\mathbf{y}_{k}^{p}-\bar{\mathbf{y}}_{k}^{p})$, where $\mathbf{y}_{k}^{p}=[\mathbf{y}_{1k}^{p,T},\dots,\mathbf{y}_{Mk}^{p,T}]^{T}\in\mathbb{C}^{MN}$, $\bar{\mathbf{y}}_{k}^{p}=[\bar{\mathbf{y}}_{1k}^{p,T},\dots,\bar{\mathbf{y}}_{Mk}^{p,T}]\in\mathbb{C}^{MN}$, and $\mathbf{\Psi}_{k}={\mathbb{E}\{(\mathbf{y}_{k}^{p}-\bar{\mathbf{y}}_{k}^{p})(\mathbf{y}_{k}^{p}-\bar{\mathbf{y}}_{k}^{p})^{H}\}}/{\tau_{p}}=\mathrm{diag}(\mathbf{\Psi}_{1k},\dots,\mathbf{\Psi}_{Mk})\in\mathbb{C}^{MN\times MN}$. The estimation error $\tilde{\mathbf{g}}_{k}=\mathbf{g}_{k}-\hat{\mathbf{g}}_{k}$ and channel estimate $\hat{\mathbf{g}}_{k}$ are also independent random variables with $\mathbb{E}\{\hat{\mathbf{g}}_{k}|\theta_{k}\}=\bar{\mathbf{g}}_{k}$, $\mathrm{Cov}\{\hat{\mathbf{g}}_{k}|\theta_{k}\}=\hat{\mathbf{R}}_{k}=p_{k}\tau_{p}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k}=\mathrm{diag}(\hat{\mathbf{R}}_{1k},\dots,\hat{\mathbf{R}}_{Mk})$, $\mathbb{E}\{\tilde{\mathbf{g}}_{k}\}=\mathbf{0}$, and $\mathrm{Cov}\{\tilde{\mathbf{g}}_{k}\}=\mathbf{C}_{k}=\check{\mathbf{R}}_{k}-p_{k}\tau_{p}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k}=\mathrm{diag}(\mathbf{C}_{1k},\dots,\mathbf{C}_{Mk})$, and $\theta_{k}=\{\theta_{mk}:m=1,\dots,M\}$, respectively.
By collecting the received data signals from all APs, the received data signal at the CPU can be denoted as $\mathbf{y}=\sum_{k=1}^{K}{\mathbf{g}_{k}x_{k}}+\mathbf{n},$ where $\mathbf{n}\sim\mathcal{N}_{\mathbb{C}}(\mathbf{0},\sigma^{2}\mathbf{I}_{MN})$ is the additive noise. Arbitrary receive combining schemes $\mathbf{v}_{k}\in\mathbb{C}^{MN}$ can be designed at the CPU based on the channel estimates $\{\hat{\mathbf{g}}_{k}:k=1,\dots,K\}$ and channel statistics to decode the data symbol transmitted by UE $k$ as
$\hat{x}_{k}=\mathbf{v}_{k}^{H}\mathbf{y}=\mathbf{v}_{k}^{H}\mathbf{g}_{k}x_{k}+\sum_{l\neq k}^{K}{\mathbf{v}_{k}^{H}\mathbf{g}_{l}x_{l}}+\mathbf{v}_{k}^{H}\mathbf{n}.$ (4)
@@
$\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF}}\!\!=\!\!\frac{p_{k}|\mathbb{E}\{\mathbf{v}_{k}^{H}\mathbf{g}_{k}\}|^{2}}{\sum\limits_{l=1}^{K}{p_{l}\mathbb{E}\{|\mathbf{v}_{k}^{H}\mathbf{g}_{l}|^{2}\}}\!-\!p_{k}|\mathbb{E}\{\mathbf{v}_{k}^{H}\mathbf{g}_{k}\}|^{2}\!+\!\sigma^{2}\mathbb{E}\{\|\mathbf{v}_{k}\|^{2}\}}.$ (6)

### v2

@@ I Introduction
To achieve demanding requirements for wireless communication networks, multiple-input multiple-output (MIMO) technology has been regarded as the promising enabler since fourth-generation (4G) wireless communication networks [1, 2, 3]. Notably, MIMO technology has gathered a big success in the fifth-generation (5G) wireless communication networks since it can deliver higher spectral efficiency in bands that feature advantageous propagation conditions [4, 5]. However, with the further development of wireless communication, sixth-generation (6G) networks are expected to meet substantially higher performance requirements [6, 7, 8, 9] and support a diverse array of emerging application scenarios [10, 11, 12, 13, 14]. To fulfill the anticipated performance enhancements, next-generation MIMO technologies have garnered significant research attention, represented by cell-free massive MIMO (CF mMIMO) [15, 16, 17], reconfigurable intelligent surfaces (RIS) [18, 19, 20], and extremely large-scale MIMO (XL-MIMO) [21, 22, 23]. Among these technologies, CF mMIMO attracted considerable research interest due to its innovative network topology and its ability to provide uniform coverage.
@@
  • For the centralized processing scheme, we study the OBE combining scheme, called centralized OBE (C-OBE), which can maximize the SE among all centralized BE-structured combining schemes. Furthermore, we present a novel closed-form computation of the C-OBE combining scheme.
@@ II Fundamentals of CF mMIMO Networks
In this paper, we study a CF mMIMO network, where $M$ APs with $N$ antennas each serve $K$ single-antenna UEs. All APs, which are connected via fronthaul to a CPU, and UEs are arbitrarily located in a wide coverage region. We define the channel response between AP $m$ and UE $k$ as $\mathbf{g}_{mk}\in\mathbb{C}^{N}$, which remains constant in each time-frequency coherence block with $\tau_{c}$ channel uses [5]. Note that the uplink pilot transmission and uplink data transmission are investigated. Thus, in each coherence time-frequency block with $\tau_{c}$ channel uses, $\tau_{p}$ and $\tau_{c}-\tau_{p}$ channel uses are reserved for pilots and data, respectively.
@@ II-A Channel Model
We consider the spatially correlated Rician fading channel model, where both the LoS component and non-line-of-sight (NLoS) component are modelled. The channel between AP $m$ and UE $k$ can be modelled as $\mathbf{g}_{mk}=\overline{\mathbf{g}}_{mk}e^{j\theta_{mk}}+\check{\mathbf{g}}_{mk},$ where $\overline{\mathbf{g}}_{mk}\in\mathbb{C}^{N}$ is the semi-deterministic LoS component with $\theta_{mk}\sim\mathcal{U}[-\pi,\pi]$ being the random phase-shift between AP $m$ and UE $k$ as in [29] and $\check{\mathbf{g}}_{mk}\sim\mathcal{N}_{\mathbb{C}}\left(\mathbf{0},\check{\mathbf{R}}_{mk}\right)$ is the stochastic NLoS component. $\check{\mathbf{R}}_{mk}\in\mathbb{C}^{N\times N}$ denotes the spatial correlation matrix with $\beta_{mk}^{\mathrm{NLoS}}={\mathrm{tr}(\check{\mathbf{R}}_{mk})}/{N}$ being the large-scale fading coefficient for the NLoS component between AP $m$ and UE $k$. Meanwhile, we assume that $\check{\mathbf{g}}_{mk}$ for every AP-UE pair are independent and channel realizations of $\check{\mathbf{g}}_{mk}$ in different coherence blocks are independent and identically distributed.
@@ II-B Channel Estimation
We use a set of $\tau_{p}$ mutually orthogonal pilot signals, $\boldsymbol{\phi}_{1},\dots,\boldsymbol{\phi}_{\tau_{p}}$, during the channel estimation with $\left\|\boldsymbol{\phi}_{t}\right\|^{2}=\tau_{p}$. Let $t_{k}\in\{1,\dots,\tau_{p}\}$ denote the index of the pilot signal sent by UE $k$. Thus, we denote the pilot signal sent by UE $k$ as $\boldsymbol{\phi}_{t_{k}}$. Due to the limited coherence block size, more than one UE would use the same pilot signal, that is $K>\tau_{p}$. We denote the subset of UEs, applying the same pilot signal as UE $k$, as $\mathcal{P}_{k}$ including UE $k$. All UEs transmit their respective allocated pilot signals and the received pilot signal at AP $m$ can be denoted as $\mathbf{Y}_{m}^{p}=\sum_{k=1}^{K}{\sqrt{p_{k}}\mathbf{g}_{mk}\boldsymbol{\phi}_{t_{k}}^{T}}+\mathbf{N}_{m}^{p}\in\mathbb{C}^{N\times\tau_{p}}$, where $p_{k}$ is the transmit power of UE $k$ and $\mathbf{N}_{m}^{p}\in\mathbb{C}^{N\times\tau_{p}}$ is the additive noise with independent $\mathcal{N}_{\mathbb{C}}\left(0,\sigma^{2}\right)$ entries with $\sigma^{2}$ being the noise power. Based on the fundamentals in [29, 5], we derive the phase-aware MMSE estimate [^1] Note that the phase-shifts vary through each coherence block. More specifically, the phase-shifts are constant over the frequency domain but vary randomly over the time domain at the same pace as the small-scale fading. Thus, each AP can be assumed to be aware to these time-varying phase-shifts. of $\mathbf{g}_{mk}$ as
@@
$\widehat{\mathbf{g}}_{mk}=\overline{\mathbf{g}}_{mk}e^{j\theta_{mk}}+\sqrt{p_{k}}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\left(\mathbf{y}_{mk}^{p}-\overline{\mathbf{y}}_{mk}^{p}\right),$ (1)
@@
where $\mathbf{y}_{mk}^{p}=\mathbf{Y}_{m}^{p}\boldsymbol{\phi}_{t_{k}}^{*}=\sum_{l\in\mathcal{P}_{k}}{\sqrt{p_{l}}\tau_{p}\mathbf{g}_{ml}}+\mathbf{n}_{mk}^{p}$, $\overline{\mathbf{y}}_{mk}^{p}=\sum_{l\in\mathcal{P}_{k}}{\sqrt{p_{l}}\tau_{p}\overline{\mathbf{g}}_{ml}}e^{j\theta_{ml}}$, $\mathbf{n}_{mk}^{p}=\mathbf{N}_{m}^{p}\boldsymbol{\phi}_{k}^{*}\sim\mathcal{N}_{\mathbb{C}}(\mathbf{0},\tau_{p}\sigma^{2}\mathbf{I}_{N})$, and $\mathbf{\Psi}_{mk}={\mathbb{E}\{(\mathbf{y}_{mk}^{p}-\overline{\mathbf{y}}_{mk}^{p})\left(\mathbf{y}_{mk}^{p}-\overline{\mathbf{y}}_{mk}^{p}\right)^{H}\}}/{\tau_{p}}=\sum_{l\in\mathcal{P}_{k}}{p_{l}\tau_{p}\check{\mathbf{R}}_{ml}}+\sigma^{2}\mathbf{I}_{N}.$ The estimation error $\tilde{\mathbf{g}}_{mk}=\mathbf{g}_{mk}-\widehat{\mathbf{g}}_{mk}$ and estimate $\widehat{\mathbf{g}}_{mk}$ are independent random variables, where $\mathbb{E}\{\widehat{\mathbf{g}}_{mk}|\theta_{mk}\}=\overline{\mathbf{g}}_{mk}e^{j\theta_{mk}}$, $\mathrm{Cov}\{\widehat{\mathbf{g}}_{mk}|\theta_{mk}\}=\widehat{\mathbf{R}}_{mk}=p_{k}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk}$, $\mathbb{E}\{\tilde{\mathbf{g}}_{mk}\}=\mathbf{0}$, and $\mathrm{Cov}\{\tilde{\mathbf{g}}_{mk}\}=\mathbf{C}_{mk}=\check{\mathbf{R}}_{mk}-p_{k}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk}$, respectively.
Remark 1.
When the phase-shifts are not known, alternative channel estimators, such as the linear MMSE (LMMSE) channel estimator $\widehat{\mathbf{g}}_{mk}=\sqrt{p_{k}}\check{\mathbf{R}}_{mk}^{\prime}\mathbf{\Psi}_{mk}^{\prime,-1}\mathbf{y}_{mk}^{p}$ [29] and least-square (LS) channel estimator $\widehat{\mathbf{g}}_{mk}=1/\sqrt{p_{k}\tau_{p}^{2}}\mathbf{y}_{mk}^{p}$ [31], can be applied, where $\check{\mathbf{R}}_{mk}^{\prime}=\check{\mathbf{R}}_{mk}+\overline{\mathbf{g}}_{mk}\overline{\mathbf{g}}_{mk}^{H}$ and $\mathbf{\Psi}_{mk}^{\prime}=\sum_{l\in\mathcal{P}_{k}}{p_{l}\tau_{p}\check{\mathbf{R}}_{ml}^{\prime}}+\sigma^{2}\mathbf{I}_{N}$.
$\overline{\mathrm{SINR}}_{k}^{\mathrm{c},\mathrm{UatF}}=\frac{p_{k}\left|\mathrm{tr}\left(\mathbf{W}_{k}^{H}\overline{\mathbf{R}}_{k}\right)\right|^{2}}{\sum_{l=1}^{K}{p_{l}\mu_{kl}}+\sum_{l\in\mathcal{P}_{k}}{p_{l}\varepsilon_{kl}}+p_{k}\omega_{k}-p_{k}\left|\mathrm{tr}\left(\mathbf{W}_{k}^{H}\overline{\mathbf{R}}_{k}\right)\right|^{2}+\sigma^{2}\mathrm{tr}\left(\mathbf{W}_{k}\overline{\mathbf{R}}_{k}\mathbf{W}_{k}^{H}\right)}$ (8)
$\omega_{k}=\!p_{k}\tau_{p}\mathrm{tr(}\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})\mathrm{tr(}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})+p_{k}\tau_{p}\mathrm{tr(}\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{kk})\mathrm{tr(}\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})-\!\mathrm{tr(}\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{kk}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})$ (10)
$+\sum\nolimits_{m_{1}=1}^{M}{\sum\nolimits_{m_{2}=1,m_{1}\neq m_{2}}^{M}{[\mathrm{tr}(\mathbf{W}_{k,m_{1}m_{1}}^{H}\overline{\mathbf{G}}_{m_{1}m_{2}kk}\mathbf{W}_{k,m_{2}m_{2}}\overline{\mathbf{G}}_{m_{1}m_{2}kk}^{H})}}+\mathrm{tr}(\mathbf{W}_{k,m_{2}m_{1}}^{H}\overline{\mathbf{G}}_{m_{2}kk}\mathbf{W}_{k,m_{2}m_{1}}\overline{\mathbf{G}}_{m_{1}kk})]$
$+\sum\nolimits_{m_{1}=1}^{M}{\mathrm{tr}(\mathbf{W}_{k,m_{1}m_{1}}^{H}\overline{\mathbf{G}}_{m_{1}kk}\mathbf{W}_{k,m_{1}m_{1}}\overline{\mathbf{G}}_{m_{1}kk})}$
@@ III-A Centralized Processing
$\mathbf{g}_{k}=\left[\mathbf{g}_{1k}^{T},\dots,\mathbf{g}_{Mk}^{T}\right]^{T}=\overline{\mathbf{g}}_{k}+\check{\mathbf{g}}_{k}\in\mathbb{C}^{MN\times 1}$ (3)
where $\overline{\mathbf{g}}_{k}=[\overline{\mathbf{g}}_{1k}^{T}e^{j\theta_{1k}},\dots,\overline{\mathbf{g}}_{Mk}^{T}e^{j\theta_{Mk}}]^{T}$ and $\check{\mathbf{g}}_{k}=[\check{\mathbf{g}}_{1k}^{T},\dots,\check{\mathbf{g}}_{Mk}^{T}]^{T}$ denote the collective LoS and NLoS components for UE $k$, respectively. Relying on the independent characteristics for channels of different AP-UE pairs, we have $\check{\mathbf{R}}_{k}=\mathbb{E}\{\check{\mathbf{g}}_{k}\check{\mathbf{g}}_{k}^{H}\}=\mathrm{diag}(\check{\mathbf{R}}_{1k},\dots,\check{\mathbf{R}}_{Mk})\in\mathbb{C}^{MN\times MN}$. Based on the received pilot signals from all APs $\{\mathbf{Y}_{m}^{p}:m=1,\dots,M\}$, the CPU can obtain the MMSE estimate of $\mathbf{g}_{k}$ as $\widehat{\mathbf{g}}_{k}=\overline{\mathbf{g}}_{k}+\sqrt{p_{k}}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}(\mathbf{y}_{k}^{p}-\overline{\mathbf{y}}_{k}^{p})$, where $\mathbf{y}_{k}^{p}=[\mathbf{y}_{1k}^{p,T},\dots,\mathbf{y}_{Mk}^{p,T}]^{T}\in\mathbb{C}^{MN}$, $\overline{\mathbf{y}}_{k}^{p}=[\overline{\mathbf{y}}_{1k}^{p,T},\dots,\overline{\mathbf{y}}_{Mk}^{p,T}]\in\mathbb{C}^{MN}$, and $\mathbf{\Psi}_{k}={\mathbb{E}\{(\mathbf{y}_{k}^{p}-\overline{\mathbf{y}}_{k}^{p})(\mathbf{y}_{k}^{p}-\overline{\mathbf{y}}_{k}^{p})^{H}\}}/{\tau_{p}}=\mathrm{diag}(\mathbf{\Psi}_{1k},\dots,\mathbf{\Psi}_{Mk})\in\mathbb{C}^{MN\times MN}$. The estimation error $\tilde{\mathbf{g}}_{k}=\mathbf{g}_{k}-\widehat{\mathbf{g}}_{k}$ and channel estimate $\widehat{\mathbf{g}}_{k}$ are also independent random variables with $\mathbb{E}\{\widehat{\mathbf{g}}_{k}|\theta_{k}\}=\overline{\mathbf{g}}_{k}$, $\mathrm{Cov}\{\widehat{\mathbf{g}}_{k}|\theta_{k}\}=\widehat{\mathbf{R}}_{k}=p_{k}\tau_{p}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k}=\mathrm{diag}(\widehat{\mathbf{R}}_{1k},\dots,\widehat{\mathbf{R}}_{Mk})$, $\mathbb{E}\{\tilde{\mathbf{g}}_{k}\}=\mathbf{0}$, $\mathrm{Cov}\{\tilde{\mathbf{g}}_{k}\}=\mathbf{C}_{k}=\check{\mathbf{R}}_{k}-p_{k}\tau_{p}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k}=\mathrm{diag}(\mathbf{C}_{1k},\dots,\mathbf{C}_{Mk})$, and $\theta_{k}=\{\theta_{mk}:m=1,\dots,M\}$, respectively.
By collecting the received data signals from all APs, the received data signal at the CPU can be denoted as $\mathbf{y}=\sum_{k=1}^{K}{\mathbf{g}_{k}x_{k}}+\mathbf{n},$ where $\mathbf{n}\sim\mathcal{N}_{\mathbb{C}}(\mathbf{0},\sigma^{2}\mathbf{I}_{MN})$ is the additive noise. Arbitrary receive combining schemes $\mathbf{v}_{k}\in\mathbb{C}^{MN}$ can be designed at the CPU based on the channel estimates $\{\widehat{\mathbf{g}}_{k}:k=1,\dots,K\}$ and channel statistics to decode the data symbol transmitted by UE $k$ as
$\widehat{x}_{k}=\mathbf{v}_{k}^{H}\mathbf{y}=\mathbf{v}_{k}^{H}\mathbf{g}_{k}x_{k}+\sum_{l\neq k}^{K}{\mathbf{v}_{k}^{H}\mathbf{g}_{l}x_{l}}+\mathbf{v}_{k}^{H}\mathbf{n}.$ (4)
@@
$\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF}}=$ (6)
$\frac{p_{k}|\mathbb{E}\{\mathbf{v}_{k}^{H}\mathbf{g}_{k}\}|^{2}}{\sum\limits_{l=1}^{K}{p_{l}\mathbb{E}\{|\mathbf{v}_{k}^{H}\mathbf{g}_{l}|^{2}\}}-p_{k}|\mathbb{E}\{\mathbf{v}_{k}^{H}\mathbf{g}_{k}\}|^{2}\!+\!\sigma^{2}\mathbb{E}\{\|\mathbf{v}_{k}\|^{2}\}}.$
@@
$\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF}}=\frac{p_{k}|\mathbb{E}\{\widehat{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\mathbf{g}_{k}\}|^{2}}{\sum\limits_{l=1}^{K}{p_{l}\mathbb{E}\{|\widehat{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\mathbf{g}_{l}|^{2}\}}-p_{k}|\mathbb{E}\{\widehat{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\mathbf{g}_{k}\}|^{2}\!+\!\sigma^{2}\mathbb{E}\{\parallel\mathbf{W}_{k}\widehat{\mathbf{g}}_{k}\parallel^{2}\}}.$ (22)
$\mathbf{w}_{k}^{*}\!\!=\!\!\left(\sum_{l=1}^{K}{p_{l}\mathbb{E}\{(\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H})^{T}\otimes(\mathbf{g}_{l}\mathbf{g}_{l}^{H})\}}-p_{k}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\})\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\})^{H}+(\mathbb{E}\{\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H}\}^{T}\otimes\mathbf{I}_{MN})\right)^{-1}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\})$ (24)